the length of a rectangle is two more than three times its width. Its perimeter is 60 inches find its dimensions
Answer by Cromlix(4381)

Width = x
Length = 2 + 3x
Perimeter = 2*width + 2*length
60 = 2x + 2(2 + 3x)
60 = 2x + 4 + 6x
Collect like terms:
2x + 6x = 60 - 4
8x = 56
x = 7
Width = 7 inches
Length = 23 inches.
Hope this helps.
:-)